can somebody help me with this question?
write the equation of the hyperbola in standard form having the given properties
center(0,0); vertices(+_3,0); foci(+_5,0)
Answer by misscrt(37) 
 can somebody help me with this question?
write the equation of the hyperbola in standard form having the given properties
center(0,0); vertices(+_3,0); foci(+_5,0)
____________________________________________________________________________
The vertices and foci are, respectively, a and c units from the center, and b^2=c^2-a^2
This hyperbola is in the form:
(x^2/a^2)-(y^2/b^2)=1
First, find the value of b^2--->b^2=c^2-a^2
=5^2-3^2
=25-9
=16
Now, you can write the standard equation for this hyperbola.
{(x^2/a^2)-(y^2/b^2)=1}
{(x^2/9)-(y^2/16)=1}
Hope that helps.
Greetings,
misscrt
|
|
|